whats the answer to 9(-5m - 1) + 5 = -229
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)


Hi,
solving
9(-5m - 1) + 5 = -229
-45m - 9 + 5 = -229
-45m = -225
m = 5
CHECKING our Answer
9 * -26 + 5 = -229

Answer by stanbon(75887)

whats the answer to 9(-5m - 1) + 5 = -229
-9m-9+5 = -229
----
-9m -4 = -229
-9m = -225
m = 25
---------------
Cheers,
Stan H.